OpenStudy (anonymous):

Write an equation for the line with a y-intercept of 3 that is perpendicular to the line 3y=2x-4, please help, I don't get it!!

OpenStudy (alexandervonhumboldt2):

perpendilucal lines have negative reciprocal slopes.

OpenStudy (alexandervonhumboldt2):

y=-3/2+3

OpenStudy (anonymous):

ok, but how did you get that

OpenStudy (anonymous):

first we need to find the slope of the equation given because a perpendicular line will have a negative reciprocal slope. To find the slope, we will put the equation in y = mx + b form where m is the slope. 3y = 2x - 4 --- divide both sides by 3 y = 2/3x - 4/3 the slope in this equation is 2/3, but we need the negative reciprocal slope for the perpendicular line. All that means is " flip " the slope and change the sign. So the negative reciprocal of 2/3 is -3/2...see how I flipped the slope and changed the sign. In y = mx + b form, the slope is in the m position and the y intercept is in the b position....so your perpendicular equation is : y = -3/2x + 3

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

no!! it seems so clear when typed out!just couldn't seem to see it. thank you

OpenStudy (anonymous):

wait i do have a question, what happened to the -4/3

OpenStudy (anonymous):

the -4/3 was the y intercept of the original equation 3y = 2x - 4. It is not needed for this problem.